the sumof three consecutive odd integers is 27.

Answer by Earlsdon(6294)

First, assign the variables: Let x = the first odd integer, then x+2 = the next consecutive odd integer and x+4 = the third consecutive odd integer.
The sum: x + x+2 + x+4 = 27 Simplify and solve for x.
3x + 6 = 27 Subtract 6 from both sides.
3x = 21 Divide both sides by 3.
x = 7
x+2 = 9
x+4 = 11
Check: 7 + 9 + 11 = 27